OpenStudy (anonymous):

Sketch the graph: -x+3y=-6 So I have to re-arrange it then sketch the graph? How do I do that?

OpenStudy (shamil98):

slope-intercept form. y = mx + b So solve for y.

OpenStudy (anonymous):

Yes but then I am stuck here: 3y=+x-6 and then divide them by 3 and thats where I am stuck on x divided by 3=?

OpenStudy (shamil98):

3y = x - 6 divide everything by 3. y = x/3 - 2 your slope (rise/run) is 1/3 , and your y-intercept is -2.

OpenStudy (shamil98):

Plot your y-intercept and form a line, by going up 1 right 3, and get the bottom side of the line by doing the opposite going down 1 and left 3.

OpenStudy (anonymous):

Thanks!